OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 8y = 8 x + 2y = 2 Choose one answer. a. zero b. 2 c. There is no x value as there is no solution to this system d. x can be any value as there are infinitely many solutions to this system

OpenStudy (anonymous):

d. x can be any value as there are infinitely many solutions to this system\[4x+8y=8\\x+2y=2\]multiple the second raw by -4 then we got\[4x+8y=8\\-4x-8y=-8\]add the second raw to the first one\[0=0\\-4x-8y=-8\]therefore we got a function with infinitely solutions.